OpenStudy (anonymous):

Dominique tossed a plastic chip for a probability experiment. The chip was blue on one side and red on the other side. The blue side landed up 50 times. The red side landed up 150 times. Based on this experiment, what is the probability that the chip will land with the blue side up on the next toss? (1 point) 33% 25% 50% 75%

OpenStudy (perl):

this is an empirical probability problem probability blue = 50 / ( 50 + 150) = 50 / 200 = 1/4 = 25%